OpenStudy (anonymous):

Find a ⋅ b. a = <5, 2>, b = <4, 5>

jimthompson5910 (jim_thompson5910):

a ⋅ b is referring to the dot product

jimthompson5910 (jim_thompson5910):

so if a = <p,q> and b = <r,s> where p,q,r,s are scalars, then a ⋅ b = p*r + q*s

jimthompson5910 (jim_thompson5910):

In this case, a = <p,q> a = <5,2> so p = 5, q = 2 b = <r,s> b = <4, 5> so r = 4, s = 5 which means a ⋅ b = p*r + q*s a ⋅ b = 5*4 + 2*5 a ⋅ b = 20 + 10 a ⋅ b = 30

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

so basically just mulitiply them and then add

jimthompson5910 (jim_thompson5910):

yes, the corresponding coordinates, then add the products